OpenStudy (anonymous):

Simplify : (3/2x)(2/3y)

OpenStudy (lgbasallote):

\[\large (\frac{3}{2x})(\frac{2}{3y}) \implies \frac{3}{2x} \times \frac{2}{3y} \implies \frac{3 \times 2}{2x \times 3y}\]

OpenStudy (anonymous):

Would is be 6/6xy?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

but you need to \(\textbf{*simplify*}\)

OpenStudy (lgbasallote):

\[\frac{6}{6xy} \implies \frac 66 \times \frac{1}{xy}\]

OpenStudy (anonymous):

Still the same? or...

OpenStudy (lgbasallote):

what's 6/6?